# OneKE: A Dockerized Schema-Guided LLM Agent-based Knowledge Extraction System

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Knowledge extraction–obtaining knowledge from data, is a critical component for a wide range of practical systems such as Knowledge Graph (KG) construction (Chen et al., 2022), Retrieval Augmentation (RAG) (Gao et al., 2023), and domain-specific applications like scientific discovery (Dagdelen et al., 2024) and intelligence analysis (Sun et al., 2023). The last decades have witnessed the development of various knowledge extraction systems (Li et al., 2023; Wei et al., 2024; Xu et al., 2024). In particular, with the emergence of Large Language Models (LLMs), new works such as InstructUIE (Shi et al., 2024), iText2KG (Lairgi et al., 2024) and AgentRE (Wang et al., 2023) have been continuously emerged. However, previous approaches still struggle to effectively extract information from raw data following complex schemas and face challenges in debugging and correcting errors when they occur.
@@ 5. Conclusion and Future Work
In this paper, we introduce OneKE, a Dockerized Schema-Guided LLM Agent-based Knowledge Extraction System. OneKE is designed for flexible application across a spectrum of extraction tasks in real-world scenarios. It can handle source texts of varying lengths and formats (such as HTML and PDF) while demonstrating the capability to adapt to diverse task configurations, generating a broad range of output schemas tailored to specific requirements. Moreover, the integration of a self-reflection mechanism enables iterative improvement informed by external feedback, thereby enhancing both accuracy and adaptability.
Long-term Maintenance. We will maintain OneKE over the long term, adding new features and fixing bugs. OneKE is built on a modular architecture that promotes extensibility, and we plan to expand the Configure Knowledge Base by integrating additional domain-specific knowledge from fields such as science, thereby broadening its applicability across domains. To advance document data extraction and comprehension, we plan to develop methodologies for the integration and analysis of diverse chart types and content. We hope OneKE can serve as a helpful tool for researchers and engineers engaging in knowledge extraction with LLMs.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715189isbn: 979-8-4007-1331-6/2025/04
@@ 1. Introduction
Knowledge extraction–obtaining knowledge from data, is a critical component for a wide range of practical systems such as Knowledge Graph (KG) construction (Chen et al., 2022), Retrieval Augmentation (RAG) (Gao et al., 2023), and domain-specific applications like scientific discovery (Dagdelen et al., 2024) and intelligence analysis (Sun et al., 2023). The last decades have witnessed the development of various knowledge extraction systems (Li et al., 2023; Wei et al., 2024; Xu et al., 2024). In particular, with the emergence of Large Language Models (LLMs), new works such as InstructUIE (Wang et al., 2023), iText2KG (Lairgi et al., 2024) and AgentRE (Shi et al., 2024) have been continuously emerged. However, previous approaches still struggle to effectively extract information from raw data following complex schemas and face challenges in debugging and correcting errors when they occur.
@@ 5. Conclusion and Future Work
In this paper, we introduce OneKE, a Dockerized Schema-Guided LLM Agent-based Knowledge Extraction System. OneKE is designed for flexible application across a spectrum of extraction tasks in real-world scenarios. It can handle source texts of varying lengths and formats (such as HTML and PDF) while demonstrating the capability to adapt to diverse task configurations, generating a broad range of output schemas tailored to specific requirements. Moreover, the integration of a self-reflection mechanism enables iterative improvement informed by external feedback, thereby improving both accuracy and adaptability.
Long-term Maintenance. We will maintain OneKE over the long term, adding new features and fixing bugs. To advance document data extraction and comprehension, we plan to develop methodologies for the integration and analysis of diverse chart types and content. We hope OneKE can serve as a helpful tool for researchers and engineers engaging in knowledge extraction with LLMs.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 62206246, No. U23B2057, No. 62176185), the Fundamental Research Funds for the Central Universities (226-2023-00138), Yongjiang Talent Introduction Programme (2021A-156-G), CIPSC-SMP-Zhipu Large Model Cross-Disciplinary Fund, Ningbo Science and Technology Special Projects under Grant No. 2023Z212. This work was supported by Ant Group and Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph.